# Face-Human-Bench: A Comprehensive Benchmark of Face and Human Understanding for Multi-modal Assistants

### v1

@@ Appendix H Ethics Statement
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0011_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0028.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0007_2.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0052.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0098.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0016_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0017_2.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0064_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0059.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0022.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0064_2.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0067.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0013_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0066.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0007_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0054.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0013_2.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0061_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0024.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0074.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0026.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0057.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0070.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0060.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0005_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0008_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0045.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0050.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0041.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0076.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0048.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0076_1.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0011.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0001_2.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0007.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0003_3.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0001_3.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0005.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0064.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0043_2.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0043.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0017.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0021.jpg)
![[Uncaptioned image]](extracted/6106861/sample_folder/sample_0017_1.jpg)

### v2

@@ Appendix H Ethics Statement
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0011_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0028.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0007_2.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0052.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0098.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0016_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0017_2.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0064_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0059.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0022.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0064_2.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0067.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0013_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0066.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0007_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0054.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0013_2.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0061_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0024.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0074.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0026.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0057.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0070.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0060.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0005_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0008_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0045.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0050.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0041.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0076.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0048.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0076_1.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0011.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0001_2.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0007.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0003_3.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0001_3.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0005.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0064.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0043_2.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0043.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0017.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0021.jpg)
![[Uncaptioned image]](extracted/6108759/sample_folder/sample_0017_1.jpg)